Technical Staff Blog

Archives 2016

Last update on .

CIS Networking group is planning to make changes to the 10G switches, cit-cs-r14-sw-1 and cit-cs-r14-sw-2.  The switch config change requires a reboot.  They will be changing and rebooting the cit-cs-r14-sw-1 on Tuesday, June 21st at 5:30am and cit-cs-r14-sw-2 on Thursday, June 23rd at 6:00am.  The GPFS NSD nodes are connected to these switches ...

Last update on .

  • New VPN configuration files are now available at software.cs.brown.edu.
  • Old VPN configuration files will expire on July 10th.
  • A short VPN outage is scheduled for 12pm on June 22nd.

 Tstaff will be renewing the Certificates used to provide the SSL connection for our VPN service.  As such, the new .ovpn configuration files ...

Last update on .

finger, floor, anyone, and the demise of rwho

Short version: The commands anyone and floor no longer work. finger/cslookup partly works but will no longer exist after this month. Its remaining functionality has been assumed by elbow.

Long version:

A week or so prior to the Great Precarious Filesystem Shift last month, we shut down our rwho service. Rwho was the system ...

Last update on .

EMC_Isilon.png On Saturday June 4th, CIS will be physically moving their EMC Isilon, the hardware that provides file services for the campus. The move is scheduled to commence at 6am and complete around 4pm. We currently have three production services relying on the Isilon: email lists, ownCloud, and time machine backups. In order to keep our ...

Last update on .

network-2400px.png This was supposed to be a trouble free morning... just power up the switch we moved downstairs yesterday and add our redundancy back in. It seems to be a trend with this project, things not going quite as we planned...

07:14: The second distribution switch was powered on at 7am. Unfortuantely, something happened that ...

Last update on .

network-2400px.png 09:53: We are currently experiencing a widespread network outage with any site outside the CS department, including wireless. Tstaff is trying to confirm it's not an issue on our end and we have reached out to the CIS networking team to get their help debugging this.

10:06: We lost the network link ...

Last update on .

network-2400px.png On May 31st @ 7am CIS will be shutting down one of our two distribution switches and, because it is directly attached, we will be shutting down our backup firewall. Because of redundancy, we do not expect this work to cause any network outages. CIS will spend the remainder of May 31st moving the switch to ...

Last update on .

ibm-gpfs.jpg Overall, the migration of GPFS to the basement went remarkably well! We lost a total of two power supplies, both failures we could sustain through redundant hardware, and ran into a few bumps along the road, but it could have been a catastrophic failure and sure wasn't. We do have some additional cleanup that ...

Last update on .

ibm-gpfs.jpg The GPFS hardware is scheduled to be moved starting May 23rd @ 06:30am. Please check this page on the day of the move for periodic status updates...

... and so it begins...

05:45: Prep work for the move has begun. We are working to ensure that the most critical services remain active during the file ...

Last update on .

postfix.png As part of our ongoing virtualization efforts, we rolled out an upgraded version of our mail relay server yesterday at noon. Unbeknownst to us, the default value for a configuration variable changed in a subtle way that affected delivery to some subdomains. This definitely affected delivery to some of our email lists. We put a ...

Monthly archives

Previous year

2015

Next year

2017